OpenStudy (anonymous):

CAN SOMEONE SOLVE THIS?? 2(3-X)-(1-X)>1

OpenStudy (jidnyesha):

just simplify the inequality further

OpenStudy (anonymous):

remove parentheses and combine like term on the left first

OpenStudy (anonymous):

clear or no?

OpenStudy (jidnyesha):

yeah thats right

OpenStudy (anonymous):

i cant solve it pls to understand it pls

OpenStudy (pratyush5):

For the time being consider that inequality as an equality sign. Just consider it for solving. It is not actually an equality.

OpenStudy (anonymous):

\[ 2(3-x)-(1-x)=2\times 3-2\times x-1+x\] is a first step

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

second step is \[6-2x-1+x\] and third step is to combine like terms, giving \[5-x\]

OpenStudy (anonymous):

now you have \[5-x>1\] add \(x\) to both sides, subtract \(1\) from both sides and you are finished

OpenStudy (jidnyesha):

2*(3-x)=(2*3)-(2*x) =6-2x so the equation becomes (6-2x)-(1-x)>1 now the negative sign in between the brackets multiplies to the second set -(1-x) becomes -1+x so now the equation becomes 6-2x-1+x>1

OpenStudy (pratyush5):

@satellite73 is correct. So what answer do you get @IneedHelppls100 ?